OpenStudy (anonymous):

Write out the first five terms of the sequence: a1 = 1 an = n - a(n -1)

OpenStudy (sriram):

a=1 right??

OpenStudy (anonymous):

yes

OpenStudy (sriram):

so 1,1,1,1,1

OpenStudy (sriram):

coz it is n-n+1=1 every term =1

myininaya (myininaya):

\[a_1=1\] \[a_n=n-a_{n-1}\] \[a_2=2-a_{2-1}=2-a_1=2-1=1\] right gj sriram

myininaya (myininaya):

\[a_3=3-a_{3-1}=3-a_{2}=3-1=2\] wait no next term is 2

myininaya (myininaya):

\[a_4=4-a_{4-1}=4-a_3=4-2=2\]

myininaya (myininaya):

\[a_5=5-a_{5-1}=5-a_4=5-2=3\]

myininaya (myininaya):

1,1,2,2,3

OpenStudy (anonymous):

what on earth? it is not 11111

myininaya (myininaya):

lol the first two terms are 1,1 i was saying gj to that part

OpenStudy (anonymous):

oh sorry i see 1 2-1=1 3-1=2 4-2=2 5-2=3 6-3=3 7-3=4 8-4=4 cool